Sales are expected to increase by 3 percent next year

TABLE 1

          Sales         $47,000

Current assets of $ 5,100,

Current liabilities      $ 6,200,

          Cost           44,650

Net fixed assets of $51,500

Owners Equity           50, 400

          Net Income 2,350

56,600

Owners Equ & Liab.      56,600

Sales are expected to increase by 3 percent next year. Net Income, that is, Net Profit Margin (NPM) is 5% of Sales. The firm has no long term debt and does not plan on acquiring any. The firm does not pay any dividends and all assets, short term liabilities, and costs vary directly with sales.

1. Base upon the above data in Table 1 the projected Total Asset will be]

[Work Space]

Circle A,B,C, or D

     A. $51,500

     B. $58,298

     C. $56,600

     D. Other $____

2. Based upon the above data how much additional financing (AFN) is required for next year?

   USE: (Current Total Assets/Current Sales) x Change in Sales – (Current Liabilities/Current Sales) x  

               Change in Sales – [(NPM x Projected Sales) x (1- Dividend ratio)]. Note that, no dividend is paid.

[Work space]

  

               Circle A, B, C, or D.                                                                                                          

A. -$ 912

B. -$   723

C.   $ 967

D.   $1,698

3. Assuming you have calculated a negative (less than zero) Additional Funds Needed (AFN) from the

      above data (Table 1) then, and as we discussed in class, which of the following is correct:

  1. The company is taking on too less debt

  2. The company is taking on too much debt

  3. The company should revise its discretionary sources of financing

  4. The company should revise its spontaneous sources of financing

  5. The company has a shortfall in cash

Answer the questions that follow this table – TABLE 2

Income Statement


For the Year 2019


  Sales

$28,400

  Cost of goods sold

21,200

  Depreciation

2,700

  Earnings before interest and taxes

$ 4,500

  Interest paid

850

  Taxable income

$ 3,650

  Taxes

1,400

  Net income

$ 2,250

Dividends $900


Balance Sheet


End-of-Year 2019


  Cash

$   550

  Accounts receivable

2,450

  Inventory

4,700

  Total current assets

$ 7,700

  Net fixed assets

16,900

  Total assets

$24,600

  Accounts payable

$ 2,700

  Long-term debt

9,800

  Common stock ($1 par value)

8,000

  Retained earnings

4,100

  Total Liab. & Equity

$24,600

4. Assume this firm decides to maintain a constant debt-equity ratio, what rate of growth can it maintain, assuming that no additional external financing is available Hint: Think about sustainable growth and define it)

Explain












5. Assume that the company wants to grow without leveraging, that is, no debt. What will be its growth rate (Think about internal growth rate and define it).


6. Assume the business is currently operating at maximum capacity. All costs, assets, and current liabilities vary directly with sales. The tax rate and the dividend payout ratio will remain constant. How much additional debt is required if no new equity is raised and sales are projected to increase by 5 percent?


7. Now assume the firm is currently operating at 84 percent of capacity. All costs and net working capital vary directly with sales. The tax rate, the profit margin, and the dividend payout ratio will remain constant. How much additional debt is required if no new

equity is raised and sales are projected to increase by 12 percent?




8. What is meant by Capital intensity? Give an example to fully describe.
Answer #1

1) Answer is (D) $56,753 assuming current assets increase at the same rate of 10.8% of sales.

2) Using  (Current Total Assets/Current Sales) x Change in Sales – (Current Liabilities/Current Sales) x

Change in Sales – [(NPM x Projected Sales) x (1- Dividend ratio)]. Note that, no dividend is paid

We get (56,600/47,000)*1,410-(6,200/47,000)*1,410-[(5%*48,410)*(1-0)=$-912

3) The company should revise its discretionary sources of financing

Assume this firm decides to maintain a constant debt-equity ratio, what rate of growth can it maintain, assuming that no additional external financing is available Hint: Think about sustainable growth and define it) -

Current Debt equity ratio = Long-term debt (9,800)/Common stock ($1 par value 8,000)+ Retained earnings (4,100) = 0.81

In order to maintain the same ratio the company has to grow at

SGR = (pm*(1-d)*(1+L)) / (T-(pm*(1-d)*(1+L)))
pm is the existing and target profit margin
d is the target dividend payout ratio
L is the target total debt to equity ratio
T is the ratio of total assets to sales

SGR = 7.92%(1-(900/2250)(1+0.81)/(0.866-(7.92%(1-(900/2250)(1+0.81)) = 5.47%

Sustainable growth is defined as the annual percentage of increase in sales that is consistent with a defined financial policy (target debt to equity ratio, target dividend payout ratio, target profit margin, target ratio of total assets to net sales).

Assume that the company wants to grow without leveraging, that is, no debt. What will be its growth rate (Think about internal growth rate and define it).

Assume that the company wants to grow without leveraging, that is, no debt. What will be its growth rate (Think about internal growth rate and define it).

Internal Growth Rate = (1 - Dividend Payout Ratio) × ROA

= (1-40%)*(2,250/24,600) = 5.49%

Assume the business is currently operating at maximum capacity. All costs, assets, and current liabilities vary directly with sales. The tax rate and the dividend payout ratio will remain constant. How much additional debt is required if no new equity is raised and sales are projected to increase by 5 percent?

If sales increass by 5%, below will be the net income using the percentage of sales ratio for all costs.

Current %ge of sales Expected
Sales $        28,400 $    29,820
COGS $        21,200 75% $    22,260
Depreciation $          2,700 10% $      2,835
EBIT $          4,500 16% $      4,725
Interest $              850 3% $          893
Taxable income $          3,650 $      3,833
Taxes $          1,400 38% $      1,470
Net income $          2,250 $      2,363

REquired debt can be calculated by using debt to income ratio=Net Operating Income/Interest = 9,800/2,363 = 4.15 times more.

Now assume the firm is currently operating at 84 percent of capacity. All costs and net working capital vary directly with sales. The tax rate, the profit margin, and the dividend payout ratio will remain constant. How much additional debt is required if no new equity is raised and sales are projected to increase by 12 percent?

Current %ge of sales Expected
Sales $        28,400 $                       31,808
COGS $        21,200 75% $                       23,744
Depreciation $          2,700 10% $                         3,024
EBIT $          4,500 16% $                         5,040
Interest $              850 3% $                             952
Taxable income $          3,650 $                         4,088
Taxes $          1,400 38% $                         1,568
Net income $          2,250 $                         2,520

Using debt to income ratio=Net Operating Income/Interest = 9,800/2,520= 3.89 times more.

What is meant by Capital intensity? Give an example to fully describe.

The term "capital intensive" refers to business processes or industries that require large amounts of investment to produce a good or service and thus have a high percentage of fixed assets, such as property, plant, and equipment (PP&E). Companies in capital-intensive industries are often marked by high levels of depreciation.

apital intensity ratio equals total assets divided by sales:

Capital Intensity Ratio = Total Assets/Sales
Capital intensity ratio is the raciprocal of the total assets' turnover ratio:

Capital Intensity Ratio = 1/Total Assets’\ Turnover Ratio

Example
Coca Cola Company (NYSE: KO) earned $46,542 million in financial year 2011-2012. Total assets at the end of the period were $79,974 million. PepsiCo's total asset turnover ratio for equivalent period was 0.94. Compare capital intensity of both the companies and conclude which one is more efficient using this single metric.

Solution

Coca Cola Company's capital intensity ratio
= Total Assets ÷ Sales
= $79,974M ÷ $46,542M
= 1.72

PepsiCo's capital intensity ratio
= 1 ÷ Asset Turnover
= 1 ÷ 0.94 = 1.06

PepsiCo seems to be using its assets more efficiently. It used only $1.06 dollars per $1 of revenue. Coca Cola Company on the other hand utilized $1.72 of assets to generate $1 of revenue.